OpenStudy (anonymous):

how do you convert 16.7 inches to yards

thomaster (thomaster):

1 yard = 36 inch 1 inch = \(\dfrac{1}{36}\) yards = 0.0277777778 yards So 16.7 inch = 0.0277777778 * 16.7 = 0.463888889 yards

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (jmark):

1 inch = 0.0277777778 yard Multiply the inch value with 0.0277777778 to convert inch to yard. So 16.7*0.0277777778=0.464 yards Calculate Inch to Yard online at http://goo.gl/0C2Y3C